OpenStudy (anonymous):

solve by factoring 5x^2+7x-66=0...use a comma to seperate answers

OpenStudy (mathstudent55):

(5x + 22)(x - 3) = 0 5x + 22 = 0 or x - 3 = 0 5x = -22 or x = 3 x = -22/5 or x = 3 -22/5, 3